# A Survey on Error-Bounded Lossy Compression for Scientific Datasets

### v1

@@ 5.2. Compression for Quantum Chemistry Simulations
5.3. Compression for Quantum Circuit Simulations
@@
Tensor network contraction simulators represent a quantum circuit as a tensor network, where quantum gates or states are represented as a tensor (Markov and Shi, 2008). Indices represent the index of a bitstring that a gate operates on. Contracting tensors requires multiplication of tensors and a summation. Tensor networks can require up to the same level of memory as full state vector simulations, depending on the circuit. Even with lower-memory footprint circuits, tensor networks can have tensors grow larger and larger as the contraction sequence advances, straining the memory resources of the system.
@@
Compression is an attractive solution to this memory footprint problem. With sufficiently high throughput compression, large state vectors can be stored in memory and processed in chunks, eliminating the need to read and write from storage. As is the case with other scientific applications, if lossy compression is utilized to achieve high compression ratios and high throughput, the impact on simulation results must be characterized and mitigated. For quantum circuit simulations, state vector fidelity and total energy of the circuit are two metrics integral for analysis; thus, loss introduced from compression must not greatly distort these values.
@@
Wu et al. (Wu et al., 2019) designed a Schrödinger algorithm-based full state vector simulation pipeline that integrates compression. MPI is used to parallelize the matrix multiplication required to apply a gate to a state vector. Each rank stores a set of compressed blocks that together compose a component of the overall state vector. Blocks are decompressed two at a time to perform a partial state vector update, and the resulting state vector piece is compressed for later use. Wu et al. explored multiple compressors to compress the state vector data, including SZ2.1 (A), SZ2.1 with complex type support (B), XOR leading-zero reduction coupled with bit-plane truncation and zstd (C), and reshuffling of real and imaginary parts together before performing C (D). The results indicate that solutions C and D achieve the highest compression ratios (30–90 for error bounds in the range [1E-5,1E-1]) across the four configurations as well as FPZIP and ZFP. These results are due to the spiky nature of state vectors: adjacent data points may not be good predictors of each other. When running a 61-qubit Grover’s search algorithm with this pipeline, the memory requirement drops from 32 exabytes to 768 terabytes using 4,096 nodes. In all, their compression integration can raise the number of qubits for a simulation by 2 to 16.
@@
Shah et al. (Shah et al., 2023) targeted tensor network-based quantum circuit simulation and proposed a GPU-based compression framework for these types of simulators. Since the target is tensors that exhibit spiky behavior, the authors applied preprocessing and postprocessing steps to cuSZ and cuSZx, the GPU implementations of SZ and SZx. Additionally, the cuSZx kernel was modified to integrate the pre- and postprocessing such that the impact on throughput was limited. At a high level, the pre- and postprocessing sparsify the tensor, leveraging the fact that many tensor values are close to zero and have little impact on the contraction result. The sparsification process requires efficient computation and storage of metadata structures, such as a bitmap, in order to boost the compressors’ performance. Their designs can yield up to 10 times greater compression ratio compared with cuSZ alone. When prioritizing throughput, the modified cuSZx kernel compressor can achieve 3 to 4 times improvement in compression ratio with limited impact on throughput.
5.4. Compression for Climate Research
@@
Climate researchers have developed some of the most extensive work (Baker and et al., 2016; Baker et al., 2017, 2022a) to quantify the impacts of lossy compression on their specific domain quantities of interest. In a series of papers researchers proposed four critical assessments: the SSIM of the visualization (and later the data with dSSIM (Baker et al., 2022b)), the p-value of the KS test, the Pearson correlation coefficient of determination, and the spatial relative error with corresponding thresholds to be established by asking a panel of domain experts if the data were distinguishable from the datasets (Baker et al., 2019). While the results of the assessments are correlated, they are independent—that is, passing any one test does not guarantee that passing the others.
These thresholds were later refined by the community. Work by Underwood and Bessac (Underwood et al., 2022) identified several weaknesses in the p-value of the KS test when used in this way, making the test too conservative in some cases and too liberal in others. Therefore, they proposed some alternative distance measures for the climate community to consider. Additionally, some papers have proposed less aggressive limits for the SSIM metric in particular (Klöwer et al., 2021).
[…]
In the work by Underwood and Bessac (Underwood et al., 2022), the OptZConfig package was used to find the largest compression ratio possible while meeting all of these quality requirements. This work found that compression ratios as high as 59.81 for the atmospheric datasets were possible using the most effective compressor (SZ3).
However, this work also highlighted many opportunities to further improve the customization of compressors for climate datasets (Underwood et al., 2022). For example, while climate datasets are often stored as 3D or 4D tensors, a correlation may or may not exist between the layers of data. Specialized compressors can detect when these layers are or are not correlated and use only the correlations that exist within the layers. Dimension permutation and fusion as such are implemented in CliZ (Jian et al., 2024). Additionally, periodicity and geographic consistency can help increase prediction accuracy in prediction-based compressors. Some climate datasets, like those used in Land or ICE, have many small fields and would benefit from common dictionary encoding optimizations.
5.5. Compression for Cosmology Research
@@
Modern cosmological simulations are used by researchers and scientists to investigate new fundamental astrophysics ideas, develop and evaluate new cosmological probes, assist in large-scale cosmological surveys, and investigate systematic uncertainties (Heitmann et al., 2019; Friesen et al., 2016). Historically such studies have required large computation- and storage-intensive simulations that are run on leadership supercomputers. Today’s supercomputers have evolved to heterogeneity with accelerator-based architectures, in particular GPU-based high-performance computing systems such as the Summit system (Summit supercomputer, 2020) at Oak Ridge National Laboratory. In order to adapt to this evolution, cosmological simulation codes such as Nyx (Almgren et al., 2013) (an adaptive mesh cosmological simulation code) have been designed to take advantage of GPU-based HPC systems and can be efficiently scaled to simulate trillions of particles on millions of cores (Almgren et al., 2013). These simulations often run on a static number of ranks, usually for the same number of compute partitions; and periodically huge amounts dump raw simulation data to the storage for future post hoc analysis. With the increase in scale of such simulations, saving all the raw data generated to disk becomes impractical because of limited storage capacity and bottlenecks in the simulation due to the I/O bandwidth required to save the data to disk  (Wan et al., 2017b, a; Cappello et al., 2019).
@@
Research has shown that general-purpose data distortion metrics, such as peak signal-to-noise ratio (PSNR), normalized root-mean-square error, mean relative error, and mean squared error, on their own cannot satisfy the demand of quality for cosmological simulation post hoc analysis (Jin et al., 2020; Grosset et al., 2020). Additionally, approaches utilizing lossy compression for scientific datasets usually apply the same compression configuration to the entire dataset (Jin et al., 2020; Tao et al., 2017a). Yet not all partitions (regions) in the cosmological simulation have the same amount of information. Cosmologists are typically interested in the dense regions since these contain halos (clusters of particles) where galaxies are formed. Hence, the sparse regions could be compressed more aggressively than the dense ones, and such an action would not impact the analysis done by cosmologists.
@@
5.6. Compression for Topology and Visualization
Topological data analysis and visualization are essential in abstracting, summarizing, and understanding scientific data in various applications, ranging from cosmology and combustion to Earth simulations and AI. Topological feature descriptors, or simply topological descriptors, provide robust capabilities for capturing, summarizing, and comparing features in scientific data. Most lossy compressors cannot preserve topological features, thus not guaranteeing topology preservation in decompressed data. Inconsistency of topology in decompressed data could lead to misinterpretation and even wrong discoveries. Below is a review of the preservation of topological features in two aspects: scalar field topology and vector field topology, both of which require customizations of existing lossy compressors.
Scalar field topological descriptors include persistence diagrams, merge trees, contour trees, Reeb graphs, and Morse and Morse–Smale complexes. Key constituents of these descriptors include critical points (maxima, minima, and saddles) and their relationships. Earlier in 2018, Soler et al. (Soler et al., 2018) developed a method to adaptively quantize data based on a given persistent simplification threshold $\epsilon$. This method guarantees the preservation of critical point pairs with a persistence larger than $\epsilon$ yet does not enforce pointwise error control. More recently, Yan et al. (Yan et al., cess) proposed TopoSZ, which builds on top of SZ1.4 with a customized quantization scheme to allow different lower/upper bounds per point based on the segmentation induced by contour trees. TopoSZ also iteratively tests whether there are false-positive/false-negative critical points in decompressed data until convergence.
@@ 5.6. Compression for Topology and Visualization
5.7. Compression for Seismic Imaging
@@
Seismic imaging is a technique for determining the seismic properties of the Earth’s subsurface (Li and Qu, 2022). The technique is extensively utilized in earthquake imaging and resource exploration, including hydrocarbon and geothermal, by energy companies such as Saudi Aramco (Huang et al., 2023e). Among all the existing seismic imaging methods, reverse time migration (RTM) is a cutting-edge one since it can effectively analyze complex seismic structures (e.g., complex velocity focusing and steep (¿70) dips imaging), compared with traditional methods such as Kirchhoff and wave equation migration  (Farmer et al., 2009).
@@
A notable limitation of RTM is the massive data it generates during its execution. In general, RTM is a full two-way wave equation and can be explained as follows. Once the input data and configurations, such as the velocity model, are prepared, RTM conducts a forward propagation using the seismic waves. This phase typically involves thousands of time steps, with each step producing a single snapshot. After this phase, RTM performs a backward propagation based on the reverse order of the generated snapshots, creating the final stacking image, which represents the overall seismic structure. In real-world use cases, a 10$\times$10$\times$8 cubic kilometers geological structure may produce up to 2,800 TB of data within only a single time step (Robein, 2016). Storing such big data into peripheral devices can degrade the runtime performance drastically, which motivates error-bounded lossy compression a promising solution to reduce the memory footprint (Huang et al., 2023e; Barbosa and Coutinho, 2023).
@@
Huang et al. (Huang et al., 2023e) proposed a hybrid lossy compression method called HyZ that combines blockwise regression (BR) and an ultra-fast prediction-based compressor SZx (Yu et al., 2022) to improve the performance of RTM overall execution. Evaluation on 3,600 snapshots of the Overthrust model shows HyZ achieves a compression ratio of 12.31x and compression/decompression speeds of 10.69 GB/s and 12.45 GB/s, respectively. Integration of HyZ into an industrial parallel RTM code improves overall performance by 6.29–6.60x over the execution without compression techniques, outperforming second-tier compressors like SZ and ZFP by up to 2.23x. HyZ also demonstrates higher fidelity than BR in preserving the visualization quality of single snapshots and the final stacking image.
@@
5.8. Compression for X-ray Light Source Data
Light sources such as the Advanced Photon Source (APS) at Argonne National Laboratory and the Linac Coherent Light Source (LCLS) at SLAC National Accelerator Center produce enormous volumes of data. With the completion of the APS upgrade project and the LCLS 2 high energy projects these systems are expected to produce data at rates exceeding 1 TB/s for some experiments and beamlines. This deluge of data presents a monumental challenge to move the data within and between sites and store the data for archival purposes. In many cases, a compression ratio target of a 10×10$\times$10 × is desired for these online workflows (Underwood et al., 2023).
So far, compression for light sources has been extensively studied in the fields of ptychography and serial crystallography. As with other disciplines, the quality of the decompressed data is of the upmost importance. However, a key challenge in assessing the quality is the automation of the analysis techniques used to study light source data—in many of these domains, the evaluation of datasets is still largely a time-consuming, manual process (Underwood et al., 2023). More work to automate these workflows would accelerate the development of compressors for these applications.

### v2

@@ 6.2 Compression for Quantum Chemistry Simulations
6.3 Compression for Quantum Circuit Simulations
@@
Tensor network contraction simulators represent a quantum circuit as a tensor network, where quantum gates or states are represented as a tensor (Markov and Shi, 2008). Indices represent the index of a bitstring that a gate operates on. Contracting tensors requires multiplication of tensors and a summation. Tensor networks can require up to the same level of memory as full state vector simulations, depending on the circuit. Even with lower-memory footprint circuits, tensor networks can have tensors grow larger and larger as the contraction sequence advances, straining the memory resources.
@@
Compression is an attractive solution to this memory footprint problem. With sufficiently high throughput compression, large state vectors can be stored in memory and processed in chunks, eliminating the need to read and write from storage.
@@
Wu et al. (Wu and et al., 2019) designed a Schrödinger algorithm-based full state vector simulation pipeline that integrates compression. MPI is used to parallelize the matrix multiplication required to apply a gate to a state vector. Each rank stores a set of compressed blocks that together compose a component of the overall state vector. Blocks are decompressed two at a time to perform a partial state vector update, and the resulting state vector piece is compressed for later use. Wu et al. explored multiple compressors to compress the state vector data, including SZ2.1 (A), SZ2.1 with complex type support (B), XOR leading-zero reduction coupled with bit-plane truncation and zstd (C), and reshuffling of real and imaginary parts together before performing C (D). The results indicate that solutions C and D achieve the highest compression ratios (30–90 for error bounds in the range [1E-5,1E-1]) across the four configurations including FPZIP and ZFP. When running a 61-qubit Grover’s search algorithm with this compressor, the memory requirement drops from 32 exabytes to 768 terabytes using 4,096 nodes. In all, their method can raise the number of qubits for a simulation by 2 to 16.
@@
Shah et al. (Shah et al., 2023) targeted tensor network-based quantum circuit simulation and proposed a GPU-based compression framework for these types of simulators. Since the target is tensors that exhibit spiky behavior, the authors applied preprocessing and postprocessing steps to cuSZ and cuSZx, the GPU implementations of SZ and SZx. Additionally, the cuSZx kernel was modified to integrate the pre- and postprocessing such that the impact on throughput was limited. At a high level, the pre- and postprocessing sparsify the tensor, leveraging the fact that many tensor values are close to zero and have little impact on the contraction result. Their designs can yield up to 10 times greater compression ratio compared with cuSZ alone. When prioritizing throughput, the modified cuSZx kernel compressor can achieve 3 to 4 times improvement in compression ratio with limited impact on throughput.
6.4 Compression for Climate Research
@@
Climate researchers have developed some of the most extensive work (Baker and et al., 2016; Baker et al., 2017, 2022a; Tintó Prims et al., 2024) to quantify the impacts of lossy compression on their specific domain quantities of interest. In a series of papers researchers proposed four critical assessments: the SSIM of the visualization (and later the data with dSSIM (Baker et al., 2022b)), the p-value of the KS test, the Pearson correlation coefficient of determination, continuous ranked probability score (CRPS) (Tintó Prims et al., 2024), and the spatial relative error with corresponding thresholds to be established by asking a panel of domain experts if the data were distinguishable from the datasets (Baker et al., 2019). These thresholds were later refined by the community. Work by Underwood and Bessac (Underwood et al., 2022) identified several weaknesses in the p-value of the KS test when used in this way, making the test too conservative in some cases and too liberal in others. Therefore, they proposed some alternative distance measures for the climate community to consider. Additionally, some papers have proposed less aggressive limits for the SSIM metric in particular (Klöwer et al., 2021).
As for the compressors customized for climate datasets, two typical examples include CliZ (Jian et al., 2024) and Huang et al. (Huang and Hoefler, 2023). CliZ (Jian et al., 2024) is a kind of prediction-based compressor, which adopts multiple optimization strategies by leveraging climate data properties such as mask-map, periodicity and geographic consistency. It also features some advanced techniques such as dimension permutation and fusion, which can further help increase prediction accuracy. Some climate datasets, like those used in Land or ICE, have many small fields and would benefit from common dictionary encoding optimizations. Experiments show that CliZ outperforms SZ3, SPERR, or QoZ1.1 on climate datasets by 20%-200% in compression ratio with comparable time cost. Unlike CliZ that pursues a good balance between compression speed/throughput and quality, Huang et al.’s work (Huang and Hoefler, 2023) focuses on maximizing the compression ratio and quality for climate data. Specifically, they propose a coordinate-based neural network which is trained to overfit the data also combined with the interpolation and quantization techniques, the resulting weight parameters serve as a compressed data for the original climate/weather grid dataset. The proposed method achieves exceptionally high compression ratios, reaching up to 3000$\times$, significantly outperforming SZ3’s compression ratio of 300$\times$, while also yielding lower mean squared error (MSE). However, this improvement comes at the cost of substantially higher computational demands. Specifically, training the network for a single climate dataset required approximately 8 hours using 4 NVIDIA RTX 3090 GPUs, whereas SZ3 completed the compression in just a few minutes using 32 CPU cores.
6.5 Compression for Cosmology Research
@@
Modern cosmological simulations are used by researchers and scientists to investigate new fundamental astrophysics ideas, develop and evaluate new cosmological probes, assist in large-scale cosmological surveys, and investigate systematic uncertainties (Heitmann et al., 2019; Friesen et al., 2016). Historically such studies have required large computation- and storage-intensive simulations that are run on leadership supercomputers. In order to adapt to this evolution, cosmological simulation codes such as Nyx (Almgren et al., 2013) (an adaptive mesh cosmological simulation code) have been designed to take advantage of GPU-based HPC systems and can be efficiently scaled to simulate trillions of particles on millions of cores (Almgren et al., 2013). These simulations often periodically dump raw simulation data to the storage for future post hoc analysis. With the increase in scale of such simulations, saving all the raw data generated to disk becomes impractical because of limited storage capacity and bottlenecks in the simulation due to the I/O bandwidth required to save the data to disk  (Wan et al., 2017b, a; Cappello et al., 2019).
@@
Research has shown that general-purpose data distortion metrics, such as peak signal-to-noise ratio (PSNR), normalized root-mean-square error, mean relative error, and mean squared error, on their own cannot satisfy the demand of quality for cosmological simulation post hoc analysis (Jin et al., 2020; Grosset et al., 2020). Additionally, approaches utilizing lossy compression for scientific datasets usually apply the same compression configuration to the entire dataset (Jin et al., 2020; Tao et al., 2017a). Yet not all partitions (regions) in the cosmological simulation have the same amount of information. Cosmologists are typically interested in the dense regions since these contain halos (clusters of particles) where galaxies are formed.
@@
6.6 Compression for Topological Data Analysis
Topological data analysis is essential in abstracting, summarizing, and understanding scientific data in various applications, ranging from cosmology and combustion to Earth simulations and AI. Topological feature descriptors, or simply topological descriptors, provide robust capabilities for capturing, summarizing, and comparing features in scientific data. Most lossy compressors cannot preserve topological features, thus not guaranteeing topology preservation in decompressed data. Inconsistency of topology in decompressed data could lead to misinterpretation and even wrong discoveries. Below is a review of the preservation of topological features in error-bounded lossy compression in two aspects: scalar field topology and vector field topology, both of which require customizations of existing lossy compressors.
Scalar field topological descriptors include persistence diagrams, merge trees, contour trees, Reeb graphs, and Morse and Morse–Smale complexes. Key constituents of these descriptors include critical points (maxima, minima, and saddles) and their relationships. Earlier in 2018, Soler et al. (Soler et al., 2018) developed a method to adaptively quantize data based on a given persistent simplification threshold $\epsilon$. This method guarantees the preservation of critical point pairs with a persistence larger than $\epsilon$ yet does not enforce pointwise error control if the persistence threshold is larger than user-specified error bounds. More recently, Yan et al. (Yan et al., cess) proposed TopoSZ, which builds on top of SZ1.4 with a customized quantization scheme to allow different lower/upper bounds per point based on the segmentation induced by contour trees. TopoSZ also iteratively tests whether there are false-positive/false-negative critical points in decompressed data until convergence.
@@ 6.6 Compression for Topological Data Analysis
6.7 Compression for Multi-resolution Data
Multi-resolution methods, such as the Adaptive Mesh Refinement (AMR) technique have been widely used in scientific simulations (Zhang et al., 2019a; Stone et al., 2020; Dubey and et al., 2022). AMR aims to reduce computational expenses while preserving the accuracy of simulation outcomes. Unlike traditional uniform mesh techniques that apply consistent resolution throughout the simulation space, AMR employs a dynamic approach. It selectively increases resolution in regions of interest, thereby optimizing computational resource usage and minimizing storage requirements. However, the space saved from using multi-resolution alone is often not enough. For instance, a multi-resolution dataset with $0.5\times 1024^{3}$ mesh points at the coarse level and $0.5\times 2048^{3}$ at the fine level could yield about 1 TB of data per snapshot. Consequently, conducting five simulations with 200 snapshots would require a total disk storage of 1 PB. Simulations used in Exascale scenarios can be even larger than that, using many thousands of points per axis (Fedeli and et al., 2022), making data size reduction a timely need.
While lossy compression has the potential to significantly reduce I/O and storage costs for multi-resolution data, a gap exists between compression and multi-resolution techniques especially because of the hierarchical nature of AMR data, where the entire dataset possesses varying resolutions/levels. Data at each level is sparse because it only covers a portion of the domain. Yet, current scientific compressors exclusively support the compression of non-sparse data with uniform resolution. Consequently, AMR data must be pre-processed before compression.
Luo et al. introduced zMesh (Luo et al., 2021), which pre-processes/reorders AMR data as a 1D array across different refinement levels using z-ordering, thereby leveraging data redundancy between these levels. However, by compressing the data in a 1D array, zMesh is unable to exploit higher-dimensional compression, leading to the loss of spatial information in higher-dimensional data.
Wang et al. propose TAC (Wang et al., 2022) to leverage high-dimensional SZ compression for each refinement level of AMR data. Specifically, to handle AMR data’s hierarchical and sparse nature, TAC introduces three pre-processing strategies and adaptively applies them based on the data’s characteristics. These strategies include: (1) an optimized sparse tensor representation (OpST) for low-density AMR levels; (2) an enhanced $k$-d tree approach for medium-density AMR levels; and (3) a padding approach (GSP) for high-density AMR levels. TAC was then extended TAC to TAC+ (Wang et al., 2024b) by specifically improving SZ2’s performance on AMR data through Shared Huffman Encoding (SHE). This approach enables individual predictions for each small block while using a single shared Huffman tree for encoding, which enhances prediction accuracy and CR for the small blocks generated during the pre-processing of AMR data.
While zMesh and TAC(+) offer offline compression solutions for AMR data, they did not delve into in-situ compression, which could notably reduce the I/O cost. AMReX supports in situ AMR compression, while it compresses high-dimensional data in 1D and utilizes a small HDF5 chunk size, leading to lower CR and reduced I/O performance. To address these issues, Wang et al. propose an effective in situ lossy compression framework for AMR simulations, AMRIC (Wang et al., 2023b), which enhances I/O performance and compression quality. Unlike AMReX’s naïve in situ compression approach, AMRIC can perform 3D compression and significantly improve SZ2’s compression performance on AMR data based on TAC+. Additionally, AMRIC incorporates the HDF5 compression filter to further enhance compression, I/O performance, and usability.
6.8 Compression for Seismic Imaging
@@
Seismic imaging is a technique for determining the seismic properties of the Earth’s subsurface (Li and Qu, 2022). The technique is extensively utilized in earthquake imaging and resource exploration, including hydrocarbon and geothermal, by energy companies such as Saudi Aramco (Huang et al., 2023c). Among all the existing seismic imaging methods, reverse time migration (RTM) is a cutting-edge one since it can effectively analyze complex seismic structures (e.g., complex velocity focusing and steep (¿70) dips imaging), compared with traditional methods such as Kirchhoff and wave equation migration  (Farmer et al., 2009).
@@
A notable limitation of RTM is the massive data it generates during its execution. In general, RTM is a full two-way wave equation and can be explained as follows. Once the input data and configurations, such as the velocity model, are prepared, RTM conducts a forward propagation using the seismic waves. This phase typically involves thousands of time steps, with each step producing a single snapshot. After this phase, RTM performs a backward propagation based on the reverse order of the generated snapshots, creating the final stacking image, which represents the overall seismic structure. In real-world use cases, a 10$\times$10$\times$8 cubic kilometers geological structure may produce up to 2,800 TB of data within only a single time step (Robein, 2016). Storing such big data into peripheral devices can degrade the runtime performance drastically, which motivates error-bounded lossy compression a promising solution to reduce the memory footprint (Huang et al., 2023c; Barbosa and Coutinho, 2023).
@@
Huang et al. (Huang et al., 2023c) proposed a hybrid lossy compression method called HyZ that combines blockwise regression (BR) and an ultra-fast prediction-based compressor SZx (Yu et al., 2022) to improve the performance of RTM overall execution. Evaluation on 3,600 snapshots of the Overthrust model shows HyZ achieves a compression ratio of 12.31x and compression/decompression speeds of 10.69 GB/s and 12.45 GB/s, respectively. Integration of HyZ into an industrial parallel RTM code improves overall performance by 6.29–6.60x over the execution without compression techniques, outperforming second-tier compressors like SZ and ZFP by up to 2.23x. HyZ also demonstrates higher fidelity than BR in preserving the visualization quality of single snapshots and the final stacking image.
@@
6.9 Compression for X-ray Light Source Data
Light sources such as the Advanced Photon Source (APS) at Argonne National Laboratory and the Linac Coherent Light Source (LCLS) at SLAC National Accelerator Center produce enormous volumes of data. With the completion of the APS upgrade project and the LCLS 2 high energy projects these systems are expected to produce data at rates exceeding 1 TB/s for some experiments and beamlines. This deluge of data presents a monumental challenge to move the data within and between sites and store the data for archival purposes. In many cases, a compression ratio target of a 10×10$\times$10 × is desired for online workflows (Underwood et al., 2023b).
So far, compression for light sources has been extensively studied in the fields of ptychography and serial crystallography. As with other disciplines, the quality of the decompressed data is of the upmost importance. However, a key challenge in assessing the quality is the automation of the analysis techniques used to study light source data—in many of these domains, the evaluation of datasets is still largely a time-consuming, manual process (Underwood et al., 2023b). More work to automate these workflows would accelerate the development of compressors for these applications.